Find f∘g if f = {(1,2), (3,4), (5,4)} and g = {(2,5), (4,3)}.

f∘g = {(2, 4), (3, 4)}
f∘g = {(2, 5), (4, 3)}
f∘g = {(2, 5), (4, 3), (5, 2)}
f∘g = {(2, 4), (4, 4)}

Answer by Edwin McCravy(20055)

Find f∘g if f = {(1,2), (3,4), (5,4)} and g = {(2,5), (4,3)}.
f∘g(x) = f(g(x)), so the domain of f∘g can only contain 
elements which are elements of the domain of g, which is 
{ 2,4 }.  

g(2) = 5, so f∘g(2) = f(g(2)) = f(5) = 4, so (2,4) ∈ f∘g 

g(4) = 3, so f∘g(4) = f(g(4)) = f(3) = 4, so (4,4) ∈ f∘g 

Therefore f∘g = { (2,4), (4,4) }

Edwin